OpenStudy (anonymous):

3x − 4y = 1 2x + 3y = 8 Which is the approximate solution to the system?

OpenStudy (anonymous):

multiply first with 2 multiply second with 3 subtract second from first

OpenStudy (anonymous):

x= 35/17 y= 22/17

OpenStudy (anonymous):

thanks ^.^